# Multi-Modal Artificial Intelligence of Embryo Grading and Pregnancy Prediction in Assisted Reproductive Technology: A Review

### v1

@@ Abstract
As a global disease, infertility has always affected human beings. The development of assisted reproductive technology can effectively solve this disease. However, the traditional in vitro fertilization-embryo transfer technology still faces many challenges in improving the success rate of pregnancy, such as the subjectivity of embryo grading and the inefficiency of integrating multi-modal data. Therefore, the introduction of artificial intelligence-based technologies is particularly crucial. This article reviews the application progress of multi-modal artificial intelligence in embryo grading and pregnancy prediction based on different data modalities (including static images, time-lapse videos and structured table data) from a new perspective, and discusses the main challenges in current research, such as the complexity of multi-modal information fusion and data scarcity.
Index Terms:
@@
artificial intelligence, assisted reproductive technology, multi-modal.List of Abbreviations
| ACC | Accuracy |
| AFC | Antral Follicle Count |
| AI | Artificial Intelligence |
| AMH | Anti-Mullerian Hormone |
| ART | Assisted Reproductive Technology |
| AUC | Area Under receiver operating characteristic Curve |
| BC | BlastCoele |
| BG | Background |
| BMI | Body Mass Index |
| CNN | Convolutional Neural Network |
| EMT | EndoMetrial Thickness |
| GNN | Graph Neural Network |
| Grad-CAM | Gradient-weighted Class Activation Mapping |
| hCG | Human Chorionic Gonadotropin |
| ICM | Inner Cell Mass |
| ICSI | IntraCytoplasmic Sperm Injection |
| IVF-ET | In Vitro Fertilization-Embryo Transfer |
| PGT | Preimplantation Genetic Testing |
| RNN | Recurrent Neural Network |
| SHAP | SHapley Additive exPlanations |
| TE | TrophEctoderm |
| TLM | Time-Lapse Microscopy |
| t-SNE | t-distributed Stochastic Neighbor Embedding |
| ZP | Zona Pellucida |
I Introduction
@@
As women age, especially over the age of 35, their fertility declines and the risk of pregnancy increases. Studies have shown that up to 10-15 of couples of childbearing age are diagnosed with infertility worldwide [1]. It is of great significance for the family and the society to help infertile couples obtain their own children through modern medical technology. For families, long-term infertility often makes both husband and wife have great psychological pressure. The emergence of Assisted Reproductive Technology (ART) provides new hope for them. This can not only greatly improve their psychological state, relieve their fertility pressure, but also enhance the happiness of the family. For the society, the popularization of ART can provide more reproductive choices for families in their countries and regions. Thus, it helps them escape from the fertility dilemma, promotes the reasonable growth of the population, and improves the fertility rate to a certain extent.
@@
In Vitro Fertilization-Embryo Transfer (IVF-ET), commonly known as test-tube baby, is a key technology in the field of assisted reproduction [2]. In this technique, oocyte and sperm are fertilized outside the body to form an embryo, which is then transferred back to the woman’s uterus to achieve pregnancy. According to clinical statistics, the pregnancy success rate of a single IVF-ET transfer cycle is only 40-50, and the final live birth rate is as low as 30 [3]. There are many reasons for this result. First of all, when selecting high-quality embryos to be transferred to the female uterus in clinical practice, it mainly relies on the embryologist’s observation of the embryo image to grade the embryo quality, and the high grade is preferred to transfer. However, this method is highly subjective and error-prone, because the results of multiple assessments by the same embryologist or the same assessment by different embryologists are quite different [4]. In addition, embryo grading has a weak correlation with final pregnancy prediction, so even high-quality embryos selected by experienced embryologists may not result in pregnancy after transfer. Secondly, the key factors affecting pregnancy include not only embryo quality, but also parental fertility indicators, such as parental age, number of high-quality embryos, endometrial thickness, sperm quality and other clinical indicators [5]. Reproductive doctors need to predict the success of pregnancy according to embryo quality and fertility indicators of the couple, which once again introduces subjectivity and uncertainty. Moreover, reproductive doctors and embryologists are not the same group of people, so the coordination between different departments once again affects the outcome of pregnancy. Therefore, the success rate of IVF-ET under the traditional clinical decision-making mode is low, and patients may need to undergo multiple embryo transfers to achieve successful pregnancy, which will lead to an increase in treatment time and cost. In addition, patients who have experienced a failure of IVF-ET treatment are often unwilling to continue to receive more treatment, because the process of ART is complex and the time span is large, which causes great psychological pressure to patients [6].
The above-mentioned problems highlight the urgent need for automated Artificial Intelligence (AI) embryo grading and pregnancy prediction in practical applications. Firstly, the automated system has higher standardization in assessing the quality of embryos, which effectively eliminates the human subjective bias, thus significantly improving the accuracy and reducing the cost of embryo selection. Secondly, automated technologies enable efficient processing of large amounts of embryo data, greatly reducing the workload of embryologists and reproductive doctors, enabling them to focus on handling complex cases and in-depth communication with patients. Finally, AI can not only provide decision support, but also have a certain degree of interpretable analysis ability, which can reveal the key influencing factors behind model predictions, and AI can integrate multi-modal information of patients to achieve highly individualized treatment recommendations. The main advantages of AI in ART are shown in Fig. 1 (a).
II Related work
@@
Dalal et al. [7] systematically introduced the tasks of gamete assessment for sperm and oocytes, and summarized a variety of methods for automated analysis, which laid a foundation for the objective and standardized assessment of gamete quality. Furthermore, Deshpande et al. [8] and Merican et al. [9] focused on the morphological characteristics of embryos and summarized automated embryo grading methods, to improving the reproducibility and predictive efficiency of traditional subjective grading. In particular, Isa et al. focused on the automated analysis of blastocyst stage embryos and summarized a series of AI-based evaluation methods, including key tasks such as blastocyst quality grading and blastocyst image segmentation. From a more macro perspective, Hanassab et al. [10] comprehensively sorted out the research status and development trend of gamete assessment and embryo assessment. Given that the ultimate goal of embryo grading is to optimize transfer success, some researchers are beginning to incorporate pregnancy outcome prediction into the evaluation framework. For example, Mapstone et al. [11] and Louis et al. [12] summarized models for embryo morphology grading and pregnancy outcome prediction.
@@
In addition to the review from the perspective of specific tasks, Dimitriadis et al. [13] proposed a systematic classification method based on the embryonic development stages (including sperm stage, oocyte stage, pronuclear stage, cleavage stage and blastocyst stage), which integrated the automated tasks involved in each stage from the perspective of time. Alsaad et al. [14] focused on time-lapse imaging data and summarized relevant AI algorithms based on its unique temporal information. At the application level, Mapari et al. [15] and Alikani et al. [16] more emphasized the potential of artificial intelligence technology in the clinical scenario of assisted reproduction, especially the possibility of combining robotics and automated operating systems to realize the whole process. In order to further improve the comprehensive understanding of the application of ART, some efforts [17][18][19][20] have attempted to integrate gamete assessment (including sperm and oocytes), embryo assessment (cleavage and blastocyst stage), ploidy prediction, preimplantation genetic testing (PGT), pregnancy prediction and other tasks into the same review framework.
@@
The main tasks of the related work are summarized as shown in Fig. 1 (b). However, these works often divide tasks roughly into several categories, and do not deeply explore the specific effects of different data modalities on the suitability of task modeling and algorithms. In fact, there are many types of data involved in assisted reproductive practice, including static images, time-lapse videos, structured tabulated data, etc. Therefore, reorganizing tasks based on data sources can not only depict the problem more accurately, but also help clarify the rationality and limitations of the model design. This review work is precisely based on this motivation. Focusing on the two main tasks: embryo grading task and pregnancy prediction task in ART, we conduct a systematic sorting and comparison starting from different data modalities. The main contributions of this paper include:
@@
We conducted a systematic review of previous studies on embryo grading and pregnancy prediction, classified by data modalities, and compared the datasets used, modeling methods, and model performance.
We summarized the ART related public datasets and discussed the core challenges in current research, especially the key issues such as multi-modal information fusion and insufficient data in the ART.
Our work not only provides a comprehensive overview of current research on automated embryo grading and pregnancy prediction techniques, but also suggests the way for future research.
III Background
III-A The first generation——IVF-ET
Step 5: The selected embryo is transferred into the mother’s uterus.
Day 1: The zygote starts the first cleavage, forming a 2-cell embryo.
Day 2: The embryo continues cleavage at the 2-4 cell stage.
Day 3: The embryo continue cleavage at the 6-8 cell stage.
III-B The second generation——ICSI
III-C The third generation——PGT
IV AI in ART
@@
With the continuous development of medical imaging and machine learning technology, AI is gradually applied to the IVF-ET process, bringing new trends to reproductive medicine. In the process of IVF-ET, embryo grading and pregnancy prediction are two key steps that affect the pregnancy outcome. Traditionally, embryo grading mainly relies on manual assessment by experienced embryologists based on the morphological characteristics of embryos, while pregnancy prediction is usually based on comprehensive analysis of parental clinical characteristics (such as age, BMI, quality of sperm) and high-quality embryo image features by reproductive doctors.
@@
As a technology that can automatically learns feature representation and performs predictive analysis from large-scale and complex data, AI has shown important potential in the field of IVF-ET. Through Convolutional Neural Network (CNN), Recurrent Neural Network (RNN) and other methods, researchers can directly extract features from static embryo images, temporal embryo images, and multi-modal clinical information to effectively improve the consistency of embryo grading and the accuracy of pregnancy prediction. It should be pointed out that, as shown in the Fig. 3, the selected methods are not the same for different tasks and different data modality. Based on this, this section systematically reviews the research of AI-based embryo grading and pregnancy prediction in recent years, focusing on the current technical path, model performance, data resources, main challenges and future development trends in clinical application, so as to provide theoretical support and practical reference for the further optimization of intelligent ART.
IV-A Embryo grading
Embryo grading plays a crucial role in ART. By carefully evaluating the morphological characteristics of embryos, this process helps clinicians to select the most promising embryos for transfer, thereby significantly improving the success rate of IVF. Traditionally, embryo grading mainly relies on the manual observation of embryo images by embryologists and a comprehensive score based on a series of morphological indicators, such as cell number, cell size uniformity, integrity of the zona pellucida, and the degree of fragmentation. With the rapid development of computer vision technology, AI-based embryo grading methods have gradually become a research hotspot.

### v2

@@ Abstract
Infertility, a pressing global health concern, affects a substantial proportion of individuals worldwide. While advancements in assisted reproductive technology (ART) have offered effective interventions, conventional in vitro fertilization-embryo transfer (IVF-ET) procedures still encounter significant hurdles in enhancing pregnancy success rates. Key challenges include the inherent subjectivity in embryo grading and the inefficiency of multi-modal data integration. Against this backdrop, the adoption of AI-driven technologies has emerged as a pivotal strategy to address these issues. This article presents a comprehensive review of the progress in AI applications for embryo grading and pregnancy prediction from a novel perspective, with a specific focus on the utilization of different modal data, such as static images, time-lapse videos, and structured tabular data. The reason for this perspective is that reorganizing tasks based on data sources can not only more accurately depict the essence of the problem but also help clarify the rationality and limitations of model design. Furthermore, this review critically examines the core challenges in contemporary research, encompassing the intricacies of multi-modal feature fusion, constraints imposed by data scarcity, limitations in model generalization capabilities, and the dynamically evolving legal and regulatory frameworks. On this basis, it explicitly identifies potential avenues for future research, aiming to provide actionable guidance for advancing the application of multi-modal AI in the field of ART.
keywords:
@@
artificial intelligence, assisted reproductive technology, multi-modal1 Introduction
@@
Study has shown that up to 10-15 of couples of childbearing age are diagnosed with infertility worldwide [1]. And as women age, especially over the age of 35, their fertility declines and the risk of pregnancy increases [2]. Infertility has long been a significant challenge for both families and society. Therefore, helping infertile couples conceive through modern medical technology is of great significance for both families and society. For families, prolonged infertility often places considerable psychological pressure on both partners [3]. The emergence of Assisted Reproductive Technology (ART) offers new hope by improving psychological well-being and overall family happiness. On a societal level, the adoption of ART provides more reproductive choices for families in different regions [4].
@@
ART consists of multiple therapeutic approaches designed to address infertility. Among these, In Vitro Fertilization-Embryo Transfer (IVF-ET) has emerged as the main and most established intervention, both in clinical practice and research [5]. In this technique, oocyte and sperm are fertilized outside the body to form an embryo, which is then transferred back to the woman’s uterus to achieve pregnancy [6]. According to clinical statistics, the pregnancy success rate of a single transfer cycle of IVF-ET is only 40-50, and the final live birth rate is as low as 30 [7]. There are many reasons for this result. First, the selection of high-quality embryos for uterine transfer primarily relies on embryologists’ visual assessment of embryo images, where higher-grade embryos are prioritized. However, this method is highly subjective and error prone, because the results of multiple assessments by the same embryologist or the same assessment by different embryologists are quite different [8]. Secondly, the key factors affecting pregnancy include not only embryo quality, but also parental fertility indicators, such as parental age, number of high-quality embryos, endometrial thickness, sperm quality and other clinical indicators [9]. Reproductive doctors need to predict the success of pregnancy according to embryo quality and fertility indicators of the couple, which once again introduces subjectivity and uncertainty. Moreover, reproductive doctors and embryologists are not the same group of people, so the coordination between different departments once again affects the outcome of pregnancy. As a result, the low success rate associated with traditional clinical decision-making in IVF-ET often necessitates multiple embryo transfers to achieve a successful pregnancy. This leads to prolonged treatment timelines and higher financial burdens [10]. Moreover, patients who have experienced failed IVF-ET cycles are often reluctant to continue treatment, primarily due to the complexity and time-intensiveness of ART, which imposes significant psychological burdens on them [11].
The aforementioned challenges underscore the urgent need for automated AI-based embryo grading and pregnancy prediction systems in clinical practice. These systems deliver transformative benefits across six key dimensions. (1) Standardization: Automated systems standardize embryo quality assessment through robust algorithms, eliminating subjective bias and ensuring consistent grading across operators and settings. (2) High efficiency: AI enables rapid processing of large-scale embryo data, significantly reducing manual workload and allowing clinicians to focus on complex cases and patient interactions. (3) High accuracy: By minimizing human error, automated grading enhances the reliability of embryo grading and pregnancy prediction, improving diagnostic precision. (4) Low cost: Automation reduces reliance on repeated manual assessments, lowering operational costs and increasing the accessibility and sustainability of advanced reproductive care. (5) Personalization: AI integrates multi-modal patient data to generate tailored treatment recommendations, supporting individualized clinical decision-making and care pathways. (6) Interpretability: AI models provide interpretable insights, revealing key factors behind predictions and fostering transparency and trust in clinical applications. The main advantages of AI in ART are shown in Fig. 1.
Compared to traditional methods for embryo grading and pregnancy prediction, AI-driven approaches have demonstrated substantial advantages, leading to their widespread adoption in the field of ART [12]. While existing studies have explored various data modalities and addressed a range of specific tasks, a systematic review that comprehensively synthesizes these findings from the perspective of underlying data modalities remains absent. To address this gap, the present review adopts a novel focus on underlying data modalities, systematically summarizing and analyzing the current state of research on AI applications in embryo grading and pregnancy prediction. This helps researchers gain a comprehensive understanding of multi-modal AI in the field of assisted reproduction.
Furthermore, this review identifies key future research directions, including: (1) advanced multi-modal feature fusion techniques to integrate diverse data sources for improved predictive performance; (2) innovative strategies to address data scarcity, such as synthetic data generation; (3) the enhancement of model generalization through cross-center validation and robust algorithm design; and (4) the development of ethical frameworks to ensure transparency, fairness, and compliance with evolving regulations. Addressing these challenges is critical, as they significantly impact the accuracy, generalizability, and usability of AI-driven embryo grading and pregnancy prediction in ART. By providing a clear theoretical foundation and contextual understanding, this review aims to facilitate future advancements in the field.
| Papers | Gamete Assessment | Embryo Assessment | Gene Prediction | Pregnancy Prediction | Robot Automation | Multi-Modal |
| [13] | ✓ | ✗ | ✗ | ✗ | ✗ | ✗ |
| [14][15] | ✗ | ✓ | ✗ | ✗ | ✗ | ✗ |
| [16][17] | ✓ | ✓ | ✗ | ✗ | ✗ | ✗ |
| [18][19][20] | ✗ | ✓ | ✗ | ✓ | ✗ | ✗ |
| [21] | ✓ | ✓ | ✓ | ✓ | ✓ | ✗ |
| [22][23] | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ |
| [24] | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ |
| [25] | ✗ | ✓ | ✓ | ✗ | ✗ | ✗ |
| [26] | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ |
| [27][28] | ✗ | ✗ | ✗ | ✗ | ✓ | ✗ |
| Ours | ✗ | ✓ | ✗ | ✓ | ✗ | ✓ |
2 Related Work
@@
Dalal et al. [13] systematically introduced the tasks of gamete assessment for sperm and oocytes, and summarized a variety of methods for automated analysis, which laid a foundation for the objective and standardized assessment of gamete quality. Furthermore, Deshpande et al. [14] focused on the morphological characteristics of embryos and summarized automated embryo grading methods, aiming to enhance the reproducibility and predictive efficiency of traditional subjective grading. In particular, Isa et al. [15] focused on the automated analysis of blastocyst stage embryos and summarized a series of AI-based evaluation methods, including key tasks such as blastocyst quality grading and blastocyst segmentation. In addition, Alsaad et al. [18] focused on time-lapse imaging data and summarized relevant AI algorithms based on its unique temporal information.
From a more macro perspective, Merican et al. [16] and Hanassab et al. [17] comprehensively sorted out the research status and development trend of gamete assessment and embryo assessment. Given that the ultimate goal of embryo grading is to optimize transfer success, some researchers are beginning to incorporate pregnancy outcome prediction into the evaluation framework. For example, Mapstone et al. [19] and Louis et al. [20] summarized models for embryo morphology grading and pregnancy outcome prediction. In order to further improve the comprehensive understanding of the application of ART, some efforts [21][22][24][25][26] attempted to integrate gamete assessment (including sperm and oocytes), embryo assessment (cleavage and blastocyst stage), preimplantation genetic testing (PGT), pregnancy prediction and other tasks into the same review framework, as shown in Fig. 1. Among them, [26] made a comprehensive summary of the historical development and technical path of ART technology.
@@
In addition to the review from the perspective of specific tasks, Dimitriadis et al. [23] proposed a systematic classification method based on the embryonic development stages (including sperm stage, oocyte stage, pronuclear stage, cleavage stage and blastocyst stage), which integrated the automated tasks involved in each stage from the perspective of time. At the application level, [27][28] more emphasized the potential of artificial intelligence technology in the clinical scenario of assisted reproduction, especially the possibility of combining robotics and automated operating systems to realize the whole process.
Although prior reviews have examined the application of AI in ART, this review offers several notable advantages as shown in Table 1. Compared with [14][15][16][17][18][19][20] our review encompasses a broader range of data modalities and a more comprehensive set of tasks, thereby providing a more systematic overview of the current researches. In contrast to [13][21][22][23][24][25], while we do not cover gamete assessment or gene prediction, we focus on the core tasks of embryo grading and pregnancy prediction in ART, with greater technical depth. The exclusion of gamete assessment is justified by its technical similarity to embryo grading, as many methodologies are transferable between the two. Gene prediction is omitted due to the limited number of studies and the lack of a well-established research framework in this area. Compared with [26][27][28], our review is confined to AI technologies and does not address robotics. This scope is deliberate, as our study is positioned at the clinical decision-making level rather than the clinical operation level, enabling a more precise response to the practical needs of clinical decision support systems.
@@
Furthermore, existing reviews have not yet delved deeply into the specific impact of different data modalities on task modeling and algorithm applicability. In fact, assisted reproductive practices involve various types of data, including static images, time-lapse videos, structured tabular data, etc. Thus, reorganizing tasks based on data sources can not only more accurately depict the essence of the problem but also help clarify the rationality and limitations of model design. This review is precisely based on this core motivation, focusing on the two major tasks in ART - embryo grading and pregnancy prediction. It conducts a systematic classification and comparative analysis from different data modalities. The main contributions of this paper include:
@@
This review presents a systematic review of prior research on embryo grading and pregnancy prediction, with a classification based on data modalities. Additionally, a comparative analysis is conducted on the datasets, modeling methods, and model performance reported across the reviewed studies.
This review systematically organizes publicly available datasets in the field of ART and explores the core challenges and future research directions in depth. Key issues include multi-modal feature fusion, data scarcity, model generalization, and the improvements of ethical frameworks.
This review can provide strong support for researchers to comprehensively understand the research trends of multi-modal AI in the field of ART.
3 Method
This study adopted a systematic literature review approach to comprehensively integrate the existing research achievements of AI in the field of embryo grading and pregnancy prediction. The inclusion and exclusion process is shown in Fig. 2.
3.1 Search strategy
In terms of the selection of retrieval databases, to balance the professionalism, comprehensiveness and technical relevance of the study, the retrieval covered six core English databases. Among them, Web of Science and Scopus, as comprehensive academic databases, covered a wide range of disciplines and can cover the cross-disciplinary research of reproductive medicine and computer science, ensuring an adequate study base. PubMed, as a professional medical database, focused on the field of clinical medicine, especially enriching clinical research literature related to ART. IEEE Xplore belonged to the technical database, focusing on the application research of technologies such as AI and machine learning (ML) in medical scenarios, and facilitating the access to technology-driven study. Google Scholar and Semantic Scholar served as supplementary databases, enabling the search of preprints, conferences and other literature, while covering interdisciplinary research and reducing search blind spots.
The search time range was set from January 1, 2020 to March 31, 2025. This move aimed to include research achievements from the past five years, fully reflecting the latest progress and research hotspots in the application of AI technology in the field, and avoiding conclusion lag due to outdated literature.
In terms of search words, we adopted a combination of ”task words + technical words”. The task words included ”in vitro fertilization - embryo transfer” (IVF-ET), ”Assisted Reproductive Technology” (ART), ”embryo grading”, and pregnancy prediction. Technical words included ”Artificial Intelligence” (AI), ”Deep Learning” (DL), ”Machine Learning” (ML), ”computer vision”, and ”Convolutional Neural Network” (CNN). Each database independently conducted searches and only retained the top 20 items to focus on core and high-impact research within the field.
In the supplementary search stage, to avoid missing key study, on the basis of database search, the ”Reference List” and ”Cited Study List” of the included literature were manually traced to supplement the original research that was in line with the theme. Through the above process, a total of 969 literature were initially retrieved.
3.2 Inclusion and exclusion criteria
Studies were included if they met all of the following predefined criteria: (1) the study’s primary focus was the application of AI to either embryo quality grading or pregnancy outcome prediction within the context of ART, excluding gamete quality assessment or genetic testing research; (2) the publication type was an original research article, encompassing peer-reviewed journal articles and full papers from major international conferences, thereby excluding reviews, commentaries, editorials, and case reports; (3) the publication was in the English language to ensure methodological accuracy and prevent misinterpretation of technical details; (4) the study was strictly limited to human embryos, excluding all research on animal models; and (5) the study provided complete and transparent reporting of three key domains: dataset characteristics (e.g., sample size, data modality), AI model specifics (e.g., algorithm type, network architecture), and model performance metrics (e.g., Accuracy, AUC).
The screening process was divided into three stages: The first stage was the initial screening and deduplication. Based on the title and abstract, 249 duplicate documents were removed, leaving 720. The second stage was the re-screening of topic matching. The full text of the studies was read through, and 443 studies that were ”not related to AI” , ”not related to ART”, ”not English”, or ”not applied research” were excluded, leaving 277. The third stage was a meticulous screening and quality verification. For the remaining 277 documents, we first excluded those with ”deviated research topics” (such as gametes assessment, gene prediction, and non-human research), and then eliminated those with ”incomplete information” (no details of dataset, no details of AI methods, and no performance indicators). A total of 92 studies that met the criteria were ultimately included. Classified by research topics, among the 92 studies, 57 focused on ”the application of AI in embryo grading”, and 35 focused on ”the application of AI in pregnancy prediction”.
4 AI in ART
@@
With the continuous development of medical imaging and machine learning technology, AI is gradually applied to the IVF-ET process, bringing new trends to reproductive medicine. In the process of IVF-ET, embryo grading and pregnancy prediction are two key steps that affect the pregnancy outcome. Traditionally, embryo grading mainly relies on manual assessment by experienced embryologists based on the morphological characteristics of embryos, while pregnancy prediction is usually based on comprehensive analysis of parental clinical characteristics (such as age, BMI, quality of sperm) and high-quality embryo image features selected by reproductive doctors.
@@
As a technology that can automatically learns feature representation and performs predictive analysis from large-scale and complex data, AI has shown important potential in the field of IVF-ET. Through Convolutional Neural Network (CNN), Recurrent Neural Network (RNN) and other methods, researchers can directly extract features from static embryo images, temporal embryo images, and multi-modal clinical information to effectively improve the consistency of embryo grading and the accuracy of pregnancy prediction. It should be pointed out that, as shown in the Fig. 3, the selected methods are not the same for different tasks and different data modality. Based on this, this section systematically reviews the research of AI-based embryo grading and pregnancy prediction in recent years, focusing on the current technical path, model performance, data modalities, main challenges and future development trends in clinical application, so as to provide theoretical support and practical reference for the further optimization of intelligent ART.
4.1 Embryo grading
Embryo grading is a critical step in ART. By assessing morphological features, it enables the selection of embryos with the highest implantation potential, thereby improving IVF success rates. Traditionally, grading has relied on manual evaluation by embryologists using criteria such as cell number, size uniformity, zona pellucida integrity, and fragmentation. With advances in computer vision, AI-based grading methods are now emerging as a promising research direction.